# Out-of-Distribution Segmentation in Autonomous Driving: Problems and State of the Art

### v1

@@ 1 Introduction
The task of semantic segmentation involves assigning each pixel in an image to a predefined class. In recent years, neural networks have become the state-of-the-art approach for semantic segmentation, achieving remarkable accuracy across applications such as autonomous driving, medical imaging, and industrial inspection. However, a fundamental limitation of these models is their reliance on closed-set assumptions; they can only classify objects that were part of their training data. When these models encounter objects that were not included in the training set during real-world deployment, they misclassify them with high confidence [42].
@@
Similar to OoD segmentation, other problem formulations, such as anomaly and open-world segmentation, share similar motivations and methodologies but differ in their definitions. Anomalies refer to rare or unusual instances within the known distribution (e.g., an atypical type of vehicle), whereas OoD objects stem from entirely unknown semantic categories. While anomalies include instances which may or may not be correctly classified by a model, OoD objects represent a fundamental mismatch between the model’s learned categories and the objects encountered during deployment. Meanwhile, open-world segmentation is the task of simultaneously detecting both known and unknown objects during inference, but the lack of benchmarks labelling both inlier and unknown classes makes it difficult to distinguish from OoD segmentation. Consequently, OoD performance is typically evaluated on an external dataset, while inlier performance is assessed on the original training dataset, blurring the boundary between the two tasks. Therefore, in this survey, we do not explicitly differentiate between both tasks but acknowledge that, ideally, they should be treated separately.
@@ 2 Problem Definition, Metrics & Dataset
Method Pixel Level Component Level OS $AP$uparrow $FPR$downarrow $sIoU_{gt}$uparrow $PPV$uparrow mean $F1$uparrow M2F GM UE Other RbA [38] 95.12 0.08 53.34 59.08 57.44 UEM [39] 94.38 0.10 57.90 41.20 43.30 Mask2Anomaly [46] 93.22 0.20 55.72 75.42 68.15 UNO [9] 93.19 0.16 70.97 72.17 77.65 EAM [15] 92.87 0.52 65.85 76.50 75.58 LR [50] 92.00 0.20 62.90 81.90 78.40 UEM (generative) 88.30 0.40 42.60 59.20 50.90 PixOOD [59] 88.90 0.30 42.68 57.49 50.82 RbA 87.85 3.33 47.44 56.16 50.42 CLS [64] 87.10 0.67 44.70 53.13 51.02 DenseHybrid [14] 87.08 0.24 45.74 50.10 50.72 Maximized Entropy [5] 85.07 0.75 47.87 62.64 48.51 DaCUP [57] 81.50 1.13 37.68 60.13 46.01 ATTA [12] 76.26 2.81 43.93 37.66 36.57 FlowEneDet [16] 73.71 0.97 42.62 42.25 39.96 SynBoost  [10] 71.34 3.15 44.28 41.75 37.57 Image Resynthesis [30] 37.71 4.70 16.61 20.48 8.38 JSRNet [56] 28.09 28.86 18.55 24.46 11.02 PGN [33] 16.52 19.69 19.42 14.89 7.39 Maximum Softmax [17] 15.72 16.60 19.72 15.93 6.25 PEBAL [52] 4.98 12.68 29.91 7.55 5.54 MC Dropout [36] 4.88 50.31 5.49 5.77 1.05 Ensemble [25] 1.06 77.20 8.63 4.71 1.28 Embedding Density [2] 0.82 46.38 35.64 2.87 2.31
@@
Table 1: Quantitative Results on SMIYC-OT: Comparison of current methods on the SMIYC-OT dataset. Methods are sorted from top to bottom according to their AP. The best result for each metric is highlighted in bold, and the second best is underlined. Methods are categorized into four groups: Mask2Former-based (M2F), Uncertainty Estimation (UE), Generative Models (GM), and others
@@
Dataset: We evaluate various methods on the SegmentMeIfYouCan Obstacle Track (SMIYC-OT) and LostAndFound-NoKnown (L&F) datasets [4], both designed for segmenting OoD objects on the road that fall outside Cityscapes’ 19 predefined classes [8]. SMIYC-OT contains 327 test images with 388 OoD instances from 31 distinct obstacles captured under diverse real-world conditions, including clear weather, night, and snow. It introduces a significant domain shift compared to Cityscapes. L&F is a refined subset of the LostAndFound dataset [44], tailored for OoD segmentation, comprising 1,043 test images with 1,709 OoD instances across seven categories. Unlike SMIYC-OT, it shares Cityscapes’ domain in lighting, camera settings, and location. Both datasets provide pixel-level annotations for roads and obstacles.
@@ 3 Current State of the Art
Method Pixel Level Component Level OS $AP$uparrow $FPR$downarrow $sIoU_{gt}$uparrow $PPV$uparrow mean $F1$uparrow M2F GM UE Other PixOOD [59] 85.07 4.46 30.18 78.47 44.41 LR [50] 83.70 100.00 49.70 95.90 72.60 DaCUP [57] 81.37 7.36 38.34 67.29 51.14 UEM [39] 81.04 1.45 39.49 55.77 46.54 FlowEneDet [16] 79.75 2.92 43.79 52.83 48.05 RbA [38] 79.25 22.57 45.56 68.90 59.31 DenseHybrid [14] 78.67 2.12 46.90 52.14 52.33 Maximized Entropy [5] 77.90 9.70 45.90 63.06 49.92 JSRNet [56] 74.17 6.59 34.28 45.89 35.97 PGN [33] 69.30 9.80 50.00 44.80 45.40 Embedding Density [2] 61.70 10.36 37.75 35.21 27.55 Image Resynthesis [30] 57.08 8.82 27.16 30.69 19.17 MC Dropout [36] 36.78 35.55 17.35 34.71 12.99 Maximum Softmax [17] 30.14 33.20 14.20 62.23 10.32 Ensemble [25] 2.89 82.03 6.66 7.64 2.68
Table 2: Quantitative Results on L&F: Comparison of current methods on the L&F dataset. Methods are sorted from top to bottom according to their AP. The best result for each metric is highlighted in bold, and the second best is underlined. Methods are categorized into four groups: Mask2Former-based (M2F), Uncertainty Estimation (UE), Generative Models (GM), and others
@@
In the following section, we start by discussing outlier exposure [18], a technique used by multiple OoD detection methods to improve performance; we then discuss the methodology behind each group of methods and some of their limitations.
@@ 3.1 Outlier Exposure
In the setting where one has knowledge about potential OoD objects a network may encounter, one can improve separation between in-distribution and OoD objects using proxy OoD objects. Proxy OoD data, often referred to as known unknowns, are samples that do not belong to the in-distribution classes but resemble potential OoD objects the model might encounter in deployment. Proxy OoD data can then be used to regularize the model’s feature space by learning representations for unknown objects; this technique is referred to as outlier exposure. The assumption here is that exposure to proxy OoD during training can help the model generalize to unseen OoD objects in deployment.
@@
Instead of real negative data as proxy OoD objects, synthetically generated outliers can be obtained by sampling a jointly trained generative model [26, 40, 66]. The motivation for using synthetically generated OoD samples is to avoid potential bias from the selected proxy dataset and to generate OoD samples that are close to the training set examples yet do not belong to any in-distribution classes. For OoD segmentation, the authors of UNO [9] used outlier patches obtained by a normalizing flow model [11] that was trained along the inlier discriminative model to generate close synthetic proxy OoD data. However, this resulted in a 4.22 decrease in $AP$ and a 1.33 decrease in the mean $F1$ score compared to using samples from ADE20k on SMIYC-OT.
@@
SMIYC-OT LostAndFound-NoKnown Pixel Level Component Level Pixel Level Component Level Method $AP$uparrow $FPR$downarrow $sIoU_{gt}$uparrow $PPV$uparrow mean $F1$uparrow $AP$uparrow $FPR$downarrow $sIoU_{gt}$uparrow $PPV$uparrow mean $F1$uparrow MaxLogit 42.73 4.26 28.28 33.97 21.89 60.19 12.88 50.63 38.98 41.32 Softmax Entropy 42.73 4.27 24.26 25.35 15.04 36.87 14.86 35.71 43.92 30.37 Maximized Entropy 81.51 0.99 46.11 36.16 38.92 85.42 1.15 57.41 53.84 58.80
Table 3: Performance comparison of simple OoD detection methods on a DINOv2 encoder with a feature pyramid network decoder (83 mIoU on Cityscapes). Methods below the line used proxy OoD data from MS-COCO for outlier supervision.
@@ 3.2 Uncertainty Estimation
A neural network with a softmax output layer can be interpreted as a statistical model $f_{\theta}(y|x)$, which assigns a probability distribution over the $n$ class labels $y\in\mathcal{C}=\{y_{1},\dots,y_{n}\}$ for each pixel in the image. This distribution depends on the parameters $\theta$ and the input $x$. Some of the first approaches proposed for OoD segmentation were to use uncertainty metrics on these probability scores like: max probability [17], softmax-entropy [5], approximations over the posterior distribution of the model’s weights [36] or predictive entropy of deep ensembles [25] to estimate the OoD score of pixels.
While these uncertainty-based metrics capture different aspects of model confidence, it is important to distinguish between epistemic, aleatoric, and predictive uncertainty. Epistemic uncertainty, also known as model uncertainty, arises from the lack of knowledge about the true model parameters and can be reduced with more training data; this is typically the uncertainty one is interested in for OoD detection, as it can be reduced by collecting more data. The epistemic uncertainty is quantified as the expected information gain between the network parameters $\theta$ and the output $\mathbb{I}[Y;\theta|x,\mathcal{D}]$. On the other hand, aleatoric uncertainty represents inherent noise in the data, such as sensor noise, and cannot be reduced by collecting more data. Aleatoric uncertainty $\mathbb{E}_{p(\theta|\mathcal{D})}[\mathbb{H}[Y|x,\theta]]$, represents the expected entropy of the output distribution given fixed model parameters. Predictive uncertainty is the total uncertainty in the model’s predictions and is composed of both epistemic and aleatoric uncertainty:
@@
Training with outlier exposure reduces epistemic uncertainty by incorporating OoD examples into the learning process, effectively shifting uncertainty from epistemic to aleatoric. As a result, softmax and predictive entropy become more reliable uncertainty measures. However, this raises a fundamental question: if a model has been trained with outliers, are these samples still truly OoD? Furthermore, could excessive exposure to outliers lead to overgeneralization, reducing the model’s ability to separate inliers from outliers?
@@
Uncertainty-based methods generally rank at the bottom of the benchmark (with the exception of Maximized Entropy [5] which utilized outlier supervision and a post-processing method called Meta-Classification [48] to eliminate false-positives), this is because closed-set models are not necessarily calibrated, often leading to overconfident predictions on unseen classes [42]. In Table 3, we evaluate simple uncertainty estimation methods on a DINOv2 [43] encoder with a FPN decoder [29], which achieves an 83 mIoU on Cityscapes. We find that unnormalized uncertainty measures, such as MaxLogit [55], significantly outperform maximum softmax. Additionally, we quantify the benefits of outlier exposure and demonstrate its substantial impact on performance. However, outlier exposure led to a $\approx$ 2 mIoU drop in inlier performance. Note that meta-classification is not included in these experiments.
@@ 3.3 Generative Models
An intuitive approach for OoD detection that does not require any labelled samples is to fit a density model to the in-distribution samples and then use the inverse of the likelihood $p(x)$ as an OoD score. One of the first approaches for OoD segmentation followed this approach and used a normalizing flow [11] to approximate the density of intermediate embeddings of a DeepLabv3+ segmentation model [6]. This approach was later extended in FlowEneDet [16], where the authors use a 2D Glow-like [20] architecture trained using correctly classified pixels as positive samples and the misclassified pixels as negative samples. The proposed architecture in FlowEneDet is an extra network that can be added to any semantic segmentation network. Beyond pure generative methods, an alternative perspective views discriminative models with softmax classifiers as energy-based models for the joint distribution $p(x,y)$ through the LogSumExp scoring function on the logits [13]. This interpretation led to approaches like PEBAL [52], which improved upon simple maximum-logit scoring for OoD segmentation. More recently, hybrid approaches such as DenseHybrid [14] have combined generative modelling of the inlier data with discriminative training of negative samples via outlier exposure to enhance OoD segmentation.
@@
Another approach for using generative models for OoD segmentation involves using generative models like variational autoencoders [21] or conditional generative adversarial networks [60] for re-synthesising an image from its predicted semantic map and comparing it with the original image [30]. The discrepancies between input and generated outputs are then used as an OoD score. This approach was further expanded to integrated uncertainty estimation with image re-synthesizing to refine the OoD segmentation in SynBoost [10]. Building on the idea of detecting OoD objects based on the difference between input and generated output, JSRNeT [56] proposes a generative discriminative reconstruction module that can be added on top of a fixed semantic segmentation network. However, instead of generating a complete image, this reconstruction module learns the appearance of the road. This method was then extended in DaCUP [57], which introduced an embedding bottleneck to better model multi-modal road appearances and improve training data utilization. Additionally, a distance-based embedding scoring feature was introduced to refine anomaly detection, along with an inpainting module that filters out false positives by identifying regions reconstructable from their surroundings. These improvements were integrated into a coupling module that combines multiple anomaly estimation signals to enhance overall performance.

### v2

@@ 1 Introduction
The task of semantic segmentation involves assigning each pixel in an image to a predefined class. In recent years, neural networks have become the state-of-the-art approach for semantic segmentation, achieving remarkable accuracy across applications such as autonomous driving, medical imaging, and industrial inspection. However, a fundamental limitation of these models is their reliance on closed-set assumptions; they are only trained to recognize and classify objects from a predefined set of categories. As a result, when deployed in open-world settings, they may encounter objects or scenes outside this training distribution. In such cases, these models often produce incorrect predictions with high confidence, as they lack an explicit mechanism for identifying unknown inputs [57].
@@
Similar to OoD segmentation, other problem formulations, such as anomaly and open-world segmentation, share similar motivations and methodologies but differ in their definitions. While there is no universally accepted definition of OoD in the literature, we adopt the following distinctions in this survey to guide our discussion. Specifically, we define anomalies as rare or unusual instances that occur within a known distribution (e.g., an atypical type of vehicle), while OoD objects stem from entirely unknown semantic categories not presented during training. While anomalies include instances that may or may not be correctly classified by a model, OoD objects represent a fundamental mismatch between the learned categories of the model and the objects encountered during deployment. Meanwhile, open-world segmentation is the task of simultaneously detecting both known and unknown objects during inference. However, the absence of benchmarks that provide labels for both inlier and unknown classes makes it challenging to distinguish between open-world and OoD segmentation in practice. As a result, it has become common practice to evaluate OoD segmentation performance on an external dataset while assessing inlier performance on the original training dataset, effectively blurring the boundary between OoD and open-world segmentation. Therefore, in this survey, we do not explicitly differentiate between both tasks but acknowledge that, ideally, they should be treated separately, as OoD segmentation methods are typically designed as stand-alone models, whereas open-world segmentation methods are incorporated in a segmentation model.
@@
The remainder of this paper is structured as follows: Section 2 defines the problem formulation, evaluation metrics, and datasets. Section 3 reviews state-of-the-art methods, summarising them and briefly mentioning some of their limitations. Section 4 discusses key limitations of current benchmarks, and Section 5 outlines future research directions.
@@ 2 Problem Definition, Metrics & Dataset
Evaluation Protocol: Evaluating OoD segmentation is more complex than standard image-level OoD detection due to the dense, pixel-wise nature of the task. In particular, performance metrics can either be computed per image and then averaged across the dataset or aggregated across all pixels before computing a global metric. In this work, we adopt the latter strategy: all predictions and ground truth labels are collected across the full test set, and metrics are computed on the resulting global distribution of in-distribution and OoD pixels. This approach reduces the influence of image-level variance and reflects the overall effectiveness of the model across the entire dataset.
@@
As the numbers of TPs, FPs, and FNs depend on the detection threshold $\tau$, the average F1 score over different $\tau$ is used to give ${\overline{F1}}$ as an evaluation metric that is less affected by the detection threshold.
@@
Method Configuration Metrics Pixel Level Component Level OE M2F GM UE Other AUPRC $\uparrow$ $\text{FPR}_{95}$downarrow $sIoU_{gt}$uparrow $PPV$uparrow ${\overline{F1}}$uparrow RbA [41] 95.12 0.08 53.34 59.08 57.44 UEM [42] 94.38 0.10 57.90 41.20 43.30 Mask2Anomaly [48] 93.22 0.20 55.72 75.42 68.15 UNO [10] 93.19 0.16 70.97 72.17 77.65 EAM [19] 92.87 0.52 65.85 76.50 75.58 LR [52] 92.00 0.20 62.90 81.90 78.40 UEM (generative) 88.30 0.40 42.60 59.20 50.90 PixOOD [61] 88.90 0.30 42.68 57.49 50.82 RbA 87.85 3.33 47.44 56.16 50.42 CLS [66] 87.10 0.67 44.70 53.13 51.02 DenseHybrid [18] 87.08 0.24 45.74 50.10 50.72 Maximized Entropy [6] 85.07 0.75 47.87 62.64 48.51 DaCUP [59] 81.50 1.13 37.68 60.13 46.01 ATTA [16] 76.26 2.81 43.93 37.66 36.57 FlowEneDet [20] 73.71 0.97 42.62 42.25 39.96 SynBoost  [11] 71.34 3.15 44.28 41.75 37.57 DOoD [15] 64.30 2.60 30.30 33.10 24.10 Image Resynthesis [33] 37.71 4.70 16.61 20.48 8.38 JSRNet [58] 28.09 28.86 18.55 24.46 11.02 PGN [36] 16.52 19.69 19.42 14.89 7.39 Maximum Softmax [21] 15.72 16.60 19.72 15.93 6.25 PEBAL [54] 4.98 12.68 29.91 7.55 5.54 MC Dropout [39] 4.88 50.31 5.49 5.77 1.05 Ensemble [29] 1.06 77.20 8.63 4.71 1.28 Embedding Density [2] 0.82 46.38 35.64 2.87 2.31
@@
Table 1: Quantitative Results on SMIYC-OT: Comparison of current methods on the SMIYC-OT dataset. Methods are sorted from top to bottom according to their AP. The best result for each metric is highlighted in bold, and the second best is underlined. Methods are categorized into four groups: Mask2Former-based (M2F), Uncertainty Estimation (UE), Generative Models (GM), and others.
@@
Dataset: We evaluate the performance of various methods on the SegmentMeIfYouCan Obstacle Track (SMIYC-OT) and LostAndFound-NoKnown (L&F) datasets [5]. Both benchmarks assess the ability to segment OoD objects on the road that do not belong to any of the 19 predefined classes in the Cityscapes dataset [9]. The SMIYC-OT dataset comprises 327 test images containing a total of 388 OoD instances from 31 distinct obstacles captured under diverse real-world conditions, including clear weather, night scenes, and snowy environments. Notably, SMIYC-OT introduces a significant domain shift compared to Cityscapes. The LostAndFound-NoKnown benchmark is a filtered and refined version of the LostAndFound dataset [46], designed to better align with the OoD segmentation task. It consists of 1,043 test images with a total of 1,709 OoD instances from seven categories. Unlike SMIYC-OT, L&F shares the same domain as Cityscapes in terms of lighting, camera settings, and geographic location. Both datasets provide pixel-level annotations for the road and obstacles in the vehicle’s path.
While other OoD segmentation datasets exist, such as Fishyscapes [3] and RoadAnomaly [33], we chose not to include them in this survey for specific reasons. Fishyscapes Static includes synthetic anomalies overlaid from unrelated datasets, which resemble the process of outlier exposure and may lead to models overfitting to the synthetic nature of these objects rather than generalizing to real-world unknowns. RoadAnomaly, on the other hand, introduces a considerable domain shift from Cityscapes due to differences in camera viewpoint and height relative to the road surface, which can result in unrealistic scenarios for standard autonomous driving applications. Consequently, we focus on SMIYC-OT and L&F as they provide a more balanced and representative evaluation setting for the real-world OoD segmentation task.
@@ 3 Current State of the Art
Method Configuration Metrics Pixel Level Component Level OE M2F GM UE Other AUPRC $\uparrow$ $FPR$downarrow $sIoU_{gt}$uparrow $PPV$uparrow ${\overline{F1}}$uparrow PixOOD [61] 85.07 4.46 30.18 78.47 44.41 LR [52] 83.70 100.00 49.70 95.90 72.60 DaCUP [59] 81.37 7.36 38.34 67.29 51.14 UEM [42] 81.04 1.45 39.49 55.77 46.54 FlowEneDet [20] 79.75 2.92 43.79 52.83 48.05 DOoD [15] 79.50 3.70 27.70 40.60 30.5 RbA [41] 79.25 22.57 45.56 68.90 59.31 DenseHybrid [18] 78.67 2.12 46.90 52.14 52.33 Maximized Entropy [6] 77.90 9.70 45.90 63.06 49.92 JSRNet [58] 74.17 6.59 34.28 45.89 35.97 PGN [36] 69.30 9.80 50.00 44.80 45.40 Embedding Density [2] 61.70 10.36 37.75 35.21 27.55 Image Resynthesis [33] 57.08 8.82 27.16 30.69 19.17 MC Dropout [39] 36.78 35.55 17.35 34.71 12.99 Maximum Softmax [21] 30.14 33.20 14.20 62.23 10.32 Ensemble [29] 2.89 82.03 6.66 7.64 2.68
Table 2: Quantitative Results on L&F: Comparison of current methods on the L&F dataset. Methods are sorted from top to bottom according to their AP. The best result for each metric is highlighted in bold, and the second best is underlined. Methods are categorized into four groups: Mask2Former-based (M2F), Uncertainty Estimation (UE), Generative Models (GM), and others.
@@
In addition, we also denote if the model used any outlier data to improve OoD detection performance, a common technique referred to as outlier exposure (OE)  [22]. In the following section, we start by discussing outlier exposure, followed by the methodology behind each group of methods and some of their limitations.
@@ 3.1 Outlier Exposure
In the setting where one has knowledge about potential OoD objects a network may encounter, one can improve separation between in-distribution and OoD objects using proxy OoD objects. Proxy OoD data, often referred to as known unknowns, are samples that do not belong to the in-distribution classes but resemble potential OoD objects the model might encounter in deployment. Proxy OoD data can then be used to regularize the feature space of the model by learning representations for unknown objects; this technique is referred to as outlier exposure. The assumption here is that exposure to proxy OoD during training can help the model generalize to unseen OoD objects in deployment.
@@
[…]
Table 3: Performance comparison of simple OoD detection methods on a DINOv2 encoder with a feature pyramid network decoder (83 % mIoU on Cityscapes). Methods below the line used proxy OoD data from MS-COCO for outlier supervision.
@@ 3.2 Uncertainty Estimation
A neural network with a softmax output layer can be interpreted as a statistical model $f_{\theta}(y|x)$, which assigns a probability distribution over the $n$ class labels $y\in\mathcal{C}=\{y_{1},\dots,y_{n}\}$ for each pixel in the image. This distribution depends on the parameters $\theta$ and the input $x$. Some of the first approaches proposed for OoD segmentation were to use uncertainty metrics on these probability scores like: max probability [21], softmax-entropy [6], approximations over the posterior distribution of the weights of the model [39] or predictive entropy of deep ensembles [29] to estimate the OoD score of pixels.
While these uncertainty-based metrics capture different aspects of model confidence, it is important to distinguish between epistemic, aleatoric, and predictive uncertainty. This decomposition is not universally agreed upon across all disciplines but follows a widely accepted framework in the probabilistic modelling community, as introduced in [13]. In this framework, epistemic uncertainty, also known as model uncertainty, arises from the lack of knowledge about the true model parameters and can be reduced with more training data; this is typically the uncertainty one is interested in for OoD detection, as it can be reduced by collecting more data. The epistemic uncertainty is quantified as the expected information gain between the network parameters $\theta$ and the output $\mathbb{I}[Y;\theta|x,\mathcal{D}]$. On the other hand, aleatoric uncertainty represents inherent noise in the data, such as sensor noise, and cannot be reduced by collecting more data. Aleatoric uncertainty $\mathbb{E}_{p(\theta|\mathcal{D})}[\mathbb{H}[Y|x,\theta]]$, represents the expected entropy of the output distribution given fixed model parameters. Predictive uncertainty is the total uncertainty in the model’s predictions and is composed of both epistemic and aleatoric uncertainty:
@@
Training with outlier exposure reduces epistemic uncertainty by incorporating OoD examples into the learning process, effectively shifting uncertainty from epistemic to aleatoric. As a result, softmax and predictive entropy become more reliable uncertainty measures. However, this raises a fundamental question: if a model has been trained with outliers, are these samples still truly OoD? Furthermore, could excessive exposure to outliers lead to overgeneralization, reducing the ability of the model to separate inliers from outliers?
@@
Uncertainty-based methods generally rank at the bottom of the benchmark (with the exception of Maximized Entropy [6] which utilized outlier supervision and a post-processing method called Meta-Classification [50] to eliminate false-positives), this is because closed-set models are not necessarily calibrated, often leading to overconfident predictions on unseen classes [44]. In Table 3, we evaluate simple uncertainty estimation methods on a DINOv2 [45] encoder with a FPN decoder [32], which achieves an 83 % mIoU on Cityscapes. We find that unnormalized uncertainty measures, such as MaxLogit [57], significantly outperform maximum softmax. Additionally, we quantify the benefits of outlier exposure and demonstrate its substantial impact on performance. However, outlier exposure led to a $\approx$ 2 % mIoU drop in inlier performance. Note that meta-classification is not included in these experiments.
@@ 3.3 Generative Models
An intuitive approach for OoD detection that does not require any labelled samples is to fit a density model to the in-distribution samples and then use the inverse of the likelihood $p(x)$ as an OoD score. One of the first approaches for OoD segmentation [2] followed this approach and used a normalizing flow [12] to approximate the density of intermediate embeddings of a DeepLabv3+ segmentation model [7]. This approach was later extended in FlowEneDet [20], where the authors use a 2D Glow-like [24] architecture trained using correctly classified pixels as positive samples and the misclassified pixels as negative samples. The proposed architecture in FlowEneDet is an extra network that can be added to any semantic segmentation network. Beyond pure generative methods, an alternative perspective views discriminative models with softmax classifiers as energy-based models for the joint distribution $p(x,y)$ through the LogSumExp scoring function on the logits [17]. This interpretation led to approaches like PEBAL [54], which improved upon simple maximum-logit scoring for OoD segmentation. More recently, hybrid approaches such as DenseHybrid [18] have combined generative modelling of the inlier data with discriminative training of negative samples via outlier exposure to enhance OoD segmentation.
@@
Another approach for using generative models for OoD segmentation involves using generative models like variational autoencoders [25] or conditional generative adversarial networks [62] for re-synthesising an image from its predicted semantic map and comparing it with the original image [33]. The discrepancies between input and generated outputs are then used as an OoD score. This approach was further expanded to integrated uncertainty estimation with image re-synthesizing to refine the OoD segmentation in SynBoost [11]. Building on the idea of detecting OoD objects based on the difference between input and generated output, JSRNeT [58] proposes a generative discriminative reconstruction module that can be added on top of a fixed semantic segmentation network. However, instead of generating a complete image, this reconstruction module learns the appearance of the road. This method was then extended in DaCUP [59], which introduced an embedding bottleneck to better model multi-modal road appearances and improve training data utilization. Additionally, a distance-based embedding scoring feature was introduced to refine anomaly detection, along with an inpainting module that filters out false positives by identifying regions reconstructable from their surroundings. These improvements were integrated into a coupling module that combines multiple anomaly estimation signals to enhance overall performance. Diffusion models have also been explored for OoD segmentation in DOoD [15], where the denoising score is used to compute per-pixel OoD scores.